Use back-substitution to solve the system of linear equations. (If there is no solution, enter NO SOLUTION. If the system is dependent, set z = a and solve for x and y in terms of a.)

10x − 4y + z = 47
−y + z = 7
z = 2

(x, y, z) =
Answer #1

... (, , t) = (5, -ѕ, 2). 1 2. ( | 0 25 ск (2 S, ko) 2- th=etor to Holt th=3+(5-) - Kol s-=h Eht+Kh- k.01 N 42 ET+R - 4 ka2+